FoxxTrot:

Simplify -9/6 divided by 3/-2

FoxxTrot:

1 attachment
Bob:

divide fractions, you turn them into positives flip improper fractions to fractions

Bob:

9/6 * 2/3

Bob:

can you solve

Shadow:

First, we simplify: \[\frac{ -9 }{ 6 } \div \frac{ 3 }{ 3 } = \frac{ -3 }{ 2 }\] We divide the first fraction by 1 (3 divided by 3 is 1) and we express it in a more simpler form. You don't have to do this now, but you would have to simplify later if you didn't... Then, recognize the following rule: \[\frac{a}{ b} \div \frac{ c }{ d } = \frac{ a }{ b} \times \frac{ d }{ c}\] Thus: \[\frac{ -3 }{ 2 } \div \frac{ 3 }{ -2 } = \frac{ -3 }{ 2 } \times \frac{ -2 }{ 3 }\] Here we can do a quick shortcut, since 3 goes into 3, one time, in the numerator of the first fraction and in the denominator of the second fraction, we can simplify them to -1 and 1 respectively. We can do this similarly for the twos, creating 1 and -1 respectively. We get: \[\frac{ -1 }{ 1 } \times \frac{ -1 }{ 1 } = -1 \times -1 = 1\] Any number divided by 1 is itself, so -1 divided by 1 is -1. Then, when you multiply two negatives, you get a positive, so -1 times -1 = 1.

Shadow:

If the short cut does not make sense to you, let me just prove it works out: Going back to: \[\frac{ -3 }{2 } \times \frac{ -2 }{ 3 } = \frac{ 6 }{ 6 } = 1\] Multiply across, -3 times -2 = 6 (neg times neg = pos) and 2 times 3 is 6. So we get 6/6. Any number divided by itself is 1. Thus, we get....1.